# Local Level Dynamic Random Partition Models for Changepoint Detection

### v1

@@ Appendix D Gesture Phase Segmentation
In this Section we represent additional plots referring to the application presented in Section 5.
Figure D.10: Human Gesture data with the phases of the video on the background, as estimated by Madeo et al. (2013). See Section 5 for details.
Figure D.11: Human Gesture data. Changepoints (dashed line) identified with the DRPM. See Section 5 for details.
Figure D.12: Human Gesture data. Changepoints (dashed line) identified with the LDDP model. See Section 5 for details.
Figure D.13: Human Gesture data. Changepoints (dashed line) identified with the WDDP model. See Section 5 for details.
Figure D.14: Human Gesture data. Changepoints (dashed line) identified with the GMDDP model. See Section 5 for details.

### v2

@@ C.2 Plots
Figure C.10 displays the trace plot of the entropy $H_{t}$ of the partition at time $t$, based on a randomly selected replicate, where
$H_{t}=\sum_{j=1}^{k_{t}}\frac{n_{jt}}{n}\log\left(\frac{n_{jt}}{n}\right),$
with $k_{t}$ denoting the number of clusters at time $t$, and $n_{jt}$ the number of units assigned to the $j$th cluster at that time.
Figure C.10: trace plot of the entropy $H_{t}$ for a randomly selected replicate at a fixed time $t$. The dashed line indicates the mean entropy value. See Section 4.1 for details.
Figure C.11: Trace plots of the number of cluster for a randomly selected replicate, for times $t=\{5,47,78,100\}$. See Section 4.1 for details.
@@ Appendix D Gesture Phase Segmentation
In this Section we represent additional plots referring to the application presented in Section 5.1.
Figure D.13: Human Gesture data with the phases of the video on the background, as estimated by Madeo et al. (2013). See Section 5.1 for details.
Figure D.14: Human Gesture data. Changepoints (dashed line) identified with the DRPM. See Section 5.1 for details.
Figure D.15: Human Gesture data. Changepoints (dashed line) identified with the LDDP model. See Section 5.1 for details.
Figure D.16: Human Gesture data. Changepoints (dashed line) identified with the WDDP model. See Section 5.1 for details.
Figure D.17: Human Gesture data. Changepoints (dashed line) identified with the GMDDP model. See Section 5.1 for details.